# Statistical Response of ENSO Complexity to Initial Condition and Model Parameter Perturbations

### v5

@@ 2.2 Statistical response and the associated information measurement
It rescales the original relative entropy $\mathcal{P}$ to the interval $[0,1)$. The rescaled relative entropy $\mathcal{E}$ takes the value of 00 if and only if $\mathcal{P}=0$. It approaches $1$ when $\mathcal{P}$ becomes infinity. The rescaled relative entropy $\mathcal{E}$ remains a monotonically increasing function in terms of the difference between the two PDFs. The rescaled relative entropy $\mathcal{E}$ will be used in all the numerical results shown in this work.
@@ 2.4 Summary
The statistical quantities in this work are computed from an ensemble simulation. It is based on a Monte Carlo simulation with $3000$ ensemble members. Such an ensemble size is large enough to reproduce the strong non-Gaussian climatology PDFs for the 6-dimensional conceptual model described in Section 3. The computational cost of using $3000$ ensemble members remains low for such a conceptual model. Therefore, it provides an accurate reference solution of the statistical response using the exact formula. It further allows us to compare the result with those using the Gaussian approximations and the quad form w/ mean and variance. Note that a small ensemble size is usually sufficient to reach reasonably accurate results when applying the Gaussian approximation or the Gaussian statistics in the quadratic form. Therefore, these approximate methods can compute the statistical response using more sophisticated and higher-dimensional operational systems. The unperturbed initial value is always given by a Gaussian distribution, centered at the observational values and equipped with a tiny variance of $10^{-4}$ along each direction. Adding such a small uncertainty to the initial value facilitates the numerical calculation of the relative entropy within a very short lead time. Other advanced techniques, such as the fluctuation-dissipation theorem (Majda et al., 2005; Kubo, 1966), can be embedded into the above methods to compute the statistical response more efficiently. These methods are helpful for more complicated systems but are not adopted here since the direct Monte Carlo simulation with $3000$ ensemble members is sufficient for the 6-dimensional conceptual model. Finally, when calculating $\nabla_{\bm{\delta}}p(\mathbf{x}_{t})$ in (6), where $p(\mathbf{x}_{t})$ is the unperturbed distribution, the derivative is approximated by a second-order accurate centered finite difference. This is achieved by adding a small numerical perturbation $\bm{\epsilon}$ with $\|\bm{\epsilon}\|\ll\|\bm{\delta}\|$ to the initial conditions or model parameters, and then computing the associated PDF at time $t$. Notably, the amplitude of the numerical perturbation $\bm{\epsilon}$ is required to be small to maintain the accuracy of computing the derivative and guarantee the dominant role of the actual perturbation $\bm{\delta}$ in computing the response. Likewise, we compute another PDF at time $t$ corresponding to the perturbation $-\bm{\epsilon}$. These two PDFs are then used to approximate the derivative numerically as
@@ 3.1 Observational data sets
In most of the studies in this work, the monthly ocean temperature and current data are from the GODAS reanalysis dataset (Behringer and Xue, 2004). The thermocline depth along the equatorial Pacific is approximated from the potential temperature as the depth of the $20^{\circ}$ isotherm contour. The analysis period is 36 years, from the start of 1982 until the end of 2017. Anomalies presented in this study are calculated by removing the monthly mean climatology of the whole period. In this work, the Niño 4 ($T_{C}$) and Niño 3 ($T_{E}$) indices are the averages of the SSTa over the CP ($160^{\circ}$-$150^{\circ}$, $5^{\circ}$-$5^{\circ}$) and EP ($150^{\circ}$-$90^{\circ}$, $5^{\circ}$-$5^{\circ}$) regions, respectively. The $h_{W}$ index is the mean thermocline depth anomaly over the western Pacific region ($120^{\circ}$-$180^{\circ}$, $5^{\circ}$-$5^{\circ}$), while the $u$ index is the mean mixed-layer zonal current in the CP region. In the last subsection of the results, the ERSSTv5 data (Huang et al., 2017), which has a longer period, is utilized to compute the number of each type of ENSO event. This longer period of observed SST data allows us to compute the number of ENSO events in Table 1 while minimizing the statistical biases. The 36-year observational period defined by the GODAS dataset may cause more errors in the statistics of interest.
The daily zonal wind data is measured at 850 hPa and is taken from the NCEP-NCAR reanalysis (Kalnay et al., 1996). It is used to describe the wind bursts in the intraseasonal scale. Removing the daily mean climatology, the anomalies are averaged over the WP region to create the wind burst index. Note that the wind lies in a faster time scale than all other state variables (daily than monthly). Although a single daily value of the $\tau$ index describing the wind anomalies has a minor effect on the SST variables, the accumulated wind over time will modulate the SST variations. Jumping up to the decadal time scale, Walker circulation strength index data are also included to illustrate the modulation of the decadal variation on the interannual ENSO characters. It is defined as the sea level pressure difference over the CP/EP ($160^{\circ}$-$80^{\circ}$, $5^{\circ}$-$5^{\circ}$) and the Indian Ocean/WP ($80^{\circ}$-$160^{\circ}$, $5^{\circ}$-$5^{\circ}$) (Kang et al., 2020). The monthly zonal SST gradient between the WP and CP region highly correlates with this Walker circulation strength index (with a simultaneous Pearson correlation coefficient of around 0.85), suggesting the significance of the air-sea interactions over the equatorial Pacific. Since the latter is more directly related to the zonal advective feedback strength over the CP region, the decadal model state variable ($I$) mainly illustrates this quantity.
@@ 3.2 Definition of different types of ENSO events
The definitions of different ENSO events are based on the average SSTa during boreal winter (DJF). The CP region is defined as $160^{\circ}$-$150^{\circ}$, $5^{\circ}$-$5^{\circ}$, with the former indicating the longitude (with $180^{\circ}$ being the Prime Meridian) and the latter the latitude (with $0^{\circ}$ being the Equator), and the EP region as $150^{\circ}$-$90^{\circ}$, $5^{\circ}$-$5^{\circ}$. Using the definitions of (Kug et al., 2009), when the EP is warmer than the CP and the EP SSTa, $T_{E}$, is greater than $0.5^{\circ}$, it is classified as an EP El Niño. Based on the classification in (Wang et al., 2019), an extreme EP El Niño event corresponds to when the maximum of the EP SSTa from April to the following March is larger than $2.5^{\circ}$. Accordingly, when the CP is warmer than the EP and the CP SSTa, $T_{C}$, is larger than $0.5^{\circ}$, it is defined as a CP El Niño. Finally, when either the $T_{C}$ or $T_{E}$ anomalies are cooler than $-0.5^{\circ}$, it is defined as a La Niña event.
@@ 4.3.1 Response to initial value perturbations for different ENSO events
Figure 8 includes the cases of the 1992 mixed CP-EP event (Columns (a)–(b)) and the 1995 CP El Niño event (Column (c)). For the mixed CP-EP event, the first two rows in Column (a) show the time evolution of the mean and variance for the unperturbed and perturbed initial conditions. The initial perturbation corresponds to the direction of the strongest response of $T_{C}$ at the lead time of 4 months. With the first 5 months, the mean response of $T_{C}$ outweighs the variance response while the response in $T_{E}$ is negligible. However, at the range of the 6- to 14-month lead times, both the mean and variance responses become significant for $T_{C}$ and $T_{E}$. Note that the unperturbed system predicts a 50% possibility of a La Niña event in October 1992 (since the mean value is $-0.5^{o}$), though the true signal does not exceed this threshold. With the perturbation, the statistical forecast indicates that a La Niña event will almost surely occur. Thus, such an initial perturbation is critical since it drives the dynamics from a possibly neutral state to a La Niña event at a lead time of around one year. Column (b) shows that if the initial perturbation is imposed based on the strongest response of $T_{E}$ at the lead time of 4 months, then the perturbation has almost no influence on the statistical evolution of $T_{C}$ and the dominant component of the response in $T_{E}$ is in the mean. The comparison between these two columns implies the role of different initial perturbations in modulating the dynamics. In addition, the mean and variance responses are not always synchronized, especially for long lead times, which further indicates the importance of utilizing the entire PDF or at least considering the variance in calculating the response. Finally, by imposing the initial perturbation along the direction of the strongest response of $T_{C}$ at the lead time of 4 months, the time evolution of the statistical response of the 1995 CP El Niño in Column (c) is similar to the CP-EP event in Column (a). That is, both the mean and the variance responses are significant for lead times up to 14 months.
@@ 4.3.2 Response to model parameter perturbations
Original system 22.6$\pm$2.8 13.4$\pm$2.4 09.2$\pm$2.1 4.5$\pm$1.7 4.0$\pm$1.5 29.9$\pm$3.0 07.5$\pm$1.8
Perturbing $\beta_{E}(I)$ ()($\uparrow$)( ↑ ) 24.4$\pm$2.8 17.0$\pm$2.9 07.4$\pm$2.7 9.3$\pm$2.4 5.4$\pm$1.5 33.8$\pm$3.0 10.0$\pm$2.1
Perturbing $\beta_{C}(I)$ ()($\uparrow$)( ↑ ) 22.2$\pm$3.3 12.2$\pm$2.4 10.0$\pm$2.6 4.3$\pm$2.0 3.8$\pm$1.9 31.0$\pm$3.0 08.2$\pm$2.3
Perturbing $\zeta_{C}$ ()($\uparrow$)( ↑ ) 22.3$\pm$2.8 11.6$\pm$1.8 10.7$\pm$1.9 4.1$\pm$1.6 3.9$\pm$1.8 30.8$\pm$3.4 07.5$\pm$2.5
Perturbing $\delta_{h}$ ()($\downarrow$)( ↓ ) 22.5$\pm$3.0 13.7$\pm$2.9 08.8$\pm$2.8 4.4$\pm$2.1 4.6$\pm$1.9 29.8$\pm$4.2 07.8$\pm$2.4
Perturbing $c_{1}(t,T_{C})$ ()($\downarrow$)( ↓ ) 26.8$\pm$2.9 12.6$\pm$2.0 14.2$\pm$2.4 4.1$\pm$1.8 5.3$\pm$1.7 31.1$\pm$2.5 06.4$\pm$2.0
@@ Acknowledgements.
[…]
$[T]$ $7.5^{\circ}$ $[\tau]$ $5$ms-1
$[t]$ $2$ months $d_{\tau}$ $2$
$r$ $0.25\varrho$ $r_{C}$ $20.75b_{0}\mu\varrho/$ 2
$r_{E}$ $23r_{C}=2.25b_{0}\mu\varrho/$ 2 $\lambda$ $2/60$
@@
$\zeta_{C}$ $20.75b_{0}\mu\varrho/$ 2 $\zeta_{E}$ $20.75b_{0}\mu\varrho/$ 2
@@
$C_{u}$ $0.03\varrho$ $m$ $2$
$\sigma(I)$ $I\varrho/5$ $\beta_{E}(I)$ $0.15(1-0.2I)\sqrt{\varrho}$
$\beta_{u}(I)$ $-0.2\beta_{E}(I)$ $\beta_{h}(I)$ $-0.4\beta_{E}(I)$
@@
$\sigma_{E}$ 00 $\sigma_{I}(I)$ $\sqrt{\lambda(4-I)I}$
@@
$\mathcal{P}[p^{\delta}(\mathbf{x}_{t}),p_{L^{\prime}}^{M,\delta}(\mathbf{x}_{t})]=\{\mathcal{S}[p_{L}^{\delta}(\mathbf{x}_{t})]-\mathcal{S}[p^{\delta}(\mathbf{x}_{t})]\}+\mathcal{P}[p_{L}^{\delta}(\mathbf{x}_{t}),p_{L^{\prime}}^{M,\delta}(\mathbf{x}_{t})]=a_{0}^{M}+\bm{\alpha}^{M}_{L^{\prime}}\cdot\overline{\mathbf{E}}_{L^{\prime}}-\mathcal{S}[p^{\delta}(\mathbf{x}_{t})],$

### v6

@@ 2.2 Statistical response and the associated information measurement
It rescales the original relative entropy $\mathcal{P}$ to the interval $[0,1)$. The rescaled relative entropy $\mathcal{E}$ takes the value of 0 if and only if $\mathcal{P}=0$. It approaches $1$ when $\mathcal{P}$ becomes infinity. The rescaled relative entropy $\mathcal{E}$ remains a monotonically increasing function in terms of the difference between the two PDFs. The rescaled relative entropy $\mathcal{E}$ will be used in all the numerical results shown in this work.
@@ 2.4 Summary
The statistical quantities in this work are computed from an ensemble simulation. It is based on a Monte Carlo simulation with $3000$ ensemble members. Such an ensemble size is large enough to reproduce the strong non-Gaussian climatology PDFs for the 6-dimensional conceptual model described in Section 3. The computational cost of using $3000$ ensemble members remains low for such a conceptual model. Therefore, it provides an accurate reference solution of the statistical response using the exact formula. It further allows us to compare the result with those using the Gaussian approximations and the quad form w/ mean and variance. Note that a small ensemble size is usually sufficient to reach reasonably accurate results when applying the Gaussian approximation or the Gaussian statistics in the quadratic form. Therefore, these approximate methods can compute the statistical response using more sophisticated and higher-dimensional operational systems. The unperturbed initial value is always given by a Gaussian distribution, centered at the observational values and equipped with a tiny variance of $10^{-4}$ along each direction. Adding such a small uncertainty to the initial value facilitates the numerical calculation of the relative entropy within a very short lead time. Other advanced techniques, such as the fluctuation-dissipation theorem (Majda et al., 2005; Kubo, 1966), can be embedded into the above methods to compute the statistical response more efficiently. These methods are helpful for more complicated systems but are not adopted here since the direct Monte Carlo simulation with $3000$ ensemble members is sufficient for the 6-dimensional conceptual model. Finally, when calculating $\nabla_{\bm{\delta}}p(\mathbf{x}_{t})$ in (6), where $p(\mathbf{x}_{t})$ is the unperturbed distribution, the derivative is approximated by a second-order accurate centered finite difference. This is achieved by adding a small numerical perturbation $\bm{\epsilon}$ with $\|\bm{\epsilon}\|\ll\|\bm{\delta}\|$ to the initial conditions or model parameters, and then computing the associated PDF at time $t$. Notably, the amplitude of the numerical perturbation $\bm{\epsilon}$ is required to be small to maintain the accuracy of computing the derivative and guarantee the dominant role of the actual perturbation $\bm{\delta}$ in computing the response. Likewise, we compute another PDF at time $t$ corresponding to the perturbation $\bm{\epsilon}$. These two PDFs are then used to approximate the derivative numerically as
@@ 3.1 Observational data sets
In most of the studies in this work, the monthly ocean temperature and current data are from the GODAS reanalysis dataset (Behringer and Xue, 2004). The thermocline depth along the equatorial Pacific is approximated from the potential temperature as the depth of the $20^{\circ}$C isotherm contour. The analysis period is 36 years, from the start of 1982 until the end of 2017. Anomalies presented in this study are calculated by removing the monthly mean climatology of the whole period. In this work, the Niño 4 ($T_{C}$) and Niño 3 ($T_{E}$) indices are the averages of the SSTa over the CP ($160^{\circ}$E-$150^{\circ}$W, $5^{\circ}$S-$5^{\circ}$N) and EP ($150^{\circ}$$90^{\circ}W$, $5^{\circ}$S-$5^{\circ}$N) regions, respectively. The $h_{W}$ index is the mean thermocline depth anomaly over the western Pacific region ($120^{\circ}$E-$180^{\circ}$, $5^{\circ}$S-$5^{\circ}$N), while the $u$ index is the mean mixed-layer zonal current in the CP region. In the last subsection of the results, the ERSSTv5 data (Huang et al., 2017), which has a longer period, is utilized to compute the number of each type of ENSO event. This longer period of observed SST data allows us to compute the number of ENSO events in Table 1 while minimizing the statistical biases. The 36-year observational period defined by the GODAS dataset may cause more errors in the statistics of interest.
The daily zonal wind data is measured at 850 hPa and is taken from the NCEP-NCAR reanalysis (Kalnay et al., 1996). It is used to describe the wind bursts in the intraseasonal scale. Removing the daily mean climatology, the anomalies are averaged over the WP region to create the wind burst index. Note that the wind lies in a faster time scale than all other state variables (daily than monthly). Although a single daily value of the $\tau$ index describing the wind anomalies has a minor effect on the SST variables, the accumulated wind over time will modulate the SST variations. Jumping up to the decadal time scale, Walker circulation strength index data are also included to illustrate the modulation of the decadal variation on the interannual ENSO characters. It is defined as the sea level pressure difference over the CP/EP ($160^{\circ}$$80^{\circ}W$, $5^{\circ}$S-$5^{\circ}$N) and the Indian Ocean/WP ($80^{\circ}$$160^{\circ}E$, $5^{\circ}$S-$5^{\circ}$N) (Kang et al., 2020). The monthly zonal SST gradient between the WP and CP region highly correlates with this Walker circulation strength index (with a simultaneous Pearson correlation coefficient of around 0.85), suggesting the significance of the air-sea interactions over the equatorial Pacific. Since the latter is more directly related to the zonal advective feedback strength over the CP region, the decadal model state variable ($I$) mainly illustrates this quantity.
@@ 3.2 Definition of different types of ENSO events
The definitions of different ENSO events are based on the average SSTa during boreal winter (DJF). The CP region is defined as $160^{\circ}$E-$150^{\circ}$W, $5^{\circ}$S-$5^{\circ}$N, with the former indicating the longitude (with $180^{\circ}$ being the Prime Meridian) and the latter the latitude (with $0^{\circ}$ being the Equator), and the EP region as $150^{\circ}$$90^{\circ}W$, $5^{\circ}$S-$5^{\circ}$N. Using the definitions of (Kug et al., 2009), when the EP is warmer than the CP and the EP SSTa, $T_{E}$, is greater than $0.5^{\circ}$C, it is classified as an EP El Niño. Based on the classification in (Wang et al., 2019), an extreme EP El Niño event corresponds to when the maximum of the EP SSTa from April to the following March is larger than $2.5^{\circ}$C. Accordingly, when the CP is warmer than the EP and the CP SSTa, $T_{C}$, is larger than $0.5^{\circ}$C, it is defined as a CP El Niño. Finally, when either the $T_{C}$ or $T_{E}$ anomalies are cooler than $0.5^{\circ}$C, it is defined as a La Niña event.
@@ 4.3.1 Response to initial value perturbations for different ENSO events
Figure 8 includes the cases of the 1992 mixed CP-EP event (Columns (a)–(b)) and the 1995 CP El Niño event (Column (c)). For the mixed CP-EP event, the first two rows in Column (a) show the time evolution of the mean and variance for the unperturbed and perturbed initial conditions. The initial perturbation corresponds to the direction of the strongest response of $T_{C}$ at the lead time of 4 months. With the first 5 months, the mean response of $T_{C}$ outweighs the variance response while the response in $T_{E}$ is negligible. However, at the range of the 6- to 14-month lead times, both the mean and variance responses become significant for $T_{C}$ and $T_{E}$. Note that the unperturbed system predicts a 50% possibility of a La Niña event in October 1992 (since the mean value is $0.5^{o}$C), though the true signal does not exceed this threshold. With the perturbation, the statistical forecast indicates that a La Niña event will almost surely occur. Thus, such an initial perturbation is critical since it drives the dynamics from a possibly neutral state to a La Niña event at a lead time of around one year. Column (b) shows that if the initial perturbation is imposed based on the strongest response of $T_{E}$ at the lead time of 4 months, then the perturbation has almost no influence on the statistical evolution of $T_{C}$ and the dominant component of the response in $T_{E}$ is in the mean. The comparison between these two columns implies the role of different initial perturbations in modulating the dynamics. In addition, the mean and variance responses are not always synchronized, especially for long lead times, which further indicates the importance of utilizing the entire PDF or at least considering the variance in calculating the response. Finally, by imposing the initial perturbation along the direction of the strongest response of $T_{C}$ at the lead time of 4 months, the time evolution of the statistical response of the 1995 CP El Niño in Column (c) is similar to the CP-EP event in Column (a). That is, both the mean and the variance responses are significant for lead times up to 14 months.
@@ 4.3.2 Response to model parameter perturbations
[…]
@@ Acknowledgements.
$[u]$ 1.51.5ms-1 $[h]$ 150150m
$[T]$ $7.5^{\circ}$C $\tau]$ 55ms-1
$[t]$ 22 months $d_{\tau}$ 22
$r$ $0.25\varrho$ $r_{C}$ $0.75b_{0}\mu\varrho/2$
$r_{E}$ $3r_{C}=2.25b_{0}\mu\varrho/2$ $\lambda$ $2/60$
@@
$\zeta_{C}$ $0.75b_{0}\mu\varrho/2$ $\zeta_{E}$ $0.75b_{0}\mu\varrho/2$
@@
$C_{u}$ $0.03\varrho$ $m$ 22
$\sigma(I)$ $I\varrho/5$ $\beta_{E}(I)$ $0.15(2-0.2I)\sqrt{\varrho}$
$\beta_{u}(I)$ $0.2\beta_{E}(I)$ $\beta_{h}(I)$ $0.4\beta_{E}(I)$
@@
$\sigma_{E}$ 0 $\sigma_{I}(I)$ $\sqrt{\lambda(4-I)I}$
@@
$\hskip-6.82881pt\mathcal{P}[p^{\delta}(\mathbf{x}_{t}),p_{L^{\prime}}^{M,\delta}(\mathbf{x}_{t})]=\{\mathcal{S}[p_{L}^{\delta}(\mathbf{x}_{t})]-\mathcal{S}[p^{\delta}(\mathbf{x}_{t})]\}+\mathcal{P}[p_{L}^{\delta}(\mathbf{x}_{t}),p_{L^{\prime}}^{M,\delta}(\mathbf{x}_{t})]=a_{0}^{M}+\bm{\alpha}^{M}_{L^{\prime}}\cdot\overline{\mathbf{E}}_{L^{\prime}}-\mathcal{S}[p^{\delta}(\mathbf{x}_{t})],$